# Improved Tangential Interpolation-based Multi-input Multi-output Modal Analysis of a Full Aircraft

### v1

@@ 5 Conclusions
This work introduces the computationally improved Loewner Framework (iLF) for the extraction of modal parameters (natural frequencies, damping ratios, and mode shapes) from multi-input multi-output (MIMO) data. Numerical and experimental validations have been carried yielding the following key findings:
@@
The iLF is pioneeringly used for the modal parameters extraction from numerical and experimental systems;
@@
The iLF is found to be as precise as N4SID and the standard LF for the extraction of modal parameters from multi-input multi-output (MIMO) data on the numerical dataset. In addition, it is not affected by the presence of close-in-frequency modes as severely as LSCE is;
@@
The iLF identification from MIMO data is shown to be robust to noise on the numerical dataset;
4.
On the same numerical dataset, the iLF identification takes less time than N4SID (2 orders of magnitude) and standard LF (an order of magnitude);
5.
For the single-output multi-output and multi-input single-output experimental cases, the N4SID identifications take more than 13 hours each, while those of iLF take less than 15 s. In addition, the N4SID identification cannot characterise the modes of the systems;
6.
Experimentally, the iLF succesfully extracts the modal parameters of trainer jet aircraft from MIMO data. Notably, to the best of the Authors’ knowledge, this is the first instance in the literature that modal parameters are extracted from this very recent dataset;
7.
A guided tutorial for the iLF for modal analysis from MIMO data is shared in the Data Availability Statement, with all codes available to replicate the findings of this study.
In summary, the iLF outperforms the said state-of-the-art methods, including its conventional counterpart, in terms of accuracy and computational performance. Hence, the authors suggest its use for the extraction of modal parameters, especially for MIMO test campaigns.
@@
Conceptualisation, G.D., and M.C.; methodology, G.D., and M.C.; software, G.D.; validation, G.D. and M.C.; formal analysis, G.D.; investigation, G.D. and M.C.; resources, G.D and M.C.; data curation, G.D.; writing—original draft preparation, G.D. and M.C.; writing—review and editing, G.D. and M.C.; visualisation, G.D.; funding acquisition, G.D and M.C..
@@
The second author is supported by the Centro Nazionale per la Mobilità Sostenibile (CNMS – Sustainable Mobility Center), Spoke 7 (Connected Networks and Smart Infrastructure), Work Package 4 (Resilience of networks, structural health monitoring and asset management).
@@ Acknowledgements
The improved Loewner Framework implementation (Data file 1) used in this work is openly available from the Zenodo repository at [DOI to be reserved] and the identification results and numerical dataset (Data file 2) supporting this study are openly available from the Zenodo repository at [https://zenodo.org/records/13254981]. This project contains the following underlying data:
@@
Data file 2. Data supporting: Improved tangential interpolation-based multi-input multi-output modal analysis of complex structures

### v2

@@ 5 Conclusions
This work introduces the computationally improved Loewner Framework (iLF) for the extraction of modal parameters (natural frequencies, damping ratios, and mode shapes) from multi-input multi-output (MIMO) data. Numerical and experimental validations have been carried out, yielding the following key findings:
@@
On the numerical dataset, the iLF is found to be as precise as N4SID and the standard LF for the extraction of modal parameters from multi-input multi-output (MIMO) data;
@@
On the same numerical dataset, the iLF identification from MIMO data is shown to be reliably robust to measurement noise, computationally less time-consuming than N4SID (2 orders of magnitude) and standard LF (an order of magnitude), and not affected by the presence of close-in-frequency modes as severely as LSCE is;
@@
On the experimental dataset, the iLF successfully extracts the modal parameters of trainer jet aircraft from MIMO data, also confirming its noise robustness and computational efficiency. Notably, to the best of the Authors’ knowledge, this is the first instance in the literature that modal parameters are extracted from this very recent dataset.
In summary, the iLF outperforms the said state-of-the-art methods, including its conventional counterpart, in terms of accuracy and computational performance. Hence, the authors suggest its use for the extraction of modal parameters, especially for MIMO test campaigns. A guided tutorial for the iLF for modal analysis from MIMO data is shared in the Data Availability Statement, with all codes available to replicate the findings of this study.
Appendix A MATLAB Code Snippets of the Improvements
A.1 Code Vectorisation
Original LF MATLAB implementation without vectorisation:
LF implementation after vectorisation:
A.2 Code Decluttering
Original LF MATLAB implementation for the conversion from descriptor state-space to a continuous state-space model:
LF implementation after decluttering:
@@ A.2 Code Decluttering
Conceptualisation, G.D. and M.C.; methodology, G.D. and M.C.; software, G.D.; validation, G.D. and M.C.; formal analysis, G.D.; investigation, G.D. and M.C.; resources, G.D and M.C.; data curation, G.D.; writing—original draft preparation, G.D. and M.C.; writing—review and editing, G.D. and M.C.; visualisation, G.D.; funding acquisition, G.D and M.C..
@@
The second author is supported by the Centro Nazionale per la Mobilità Sostenibile (MOST – Sustainable Mobility Center), Spoke 7 (Cooperative Connected and Automated Mobility and Smart Infrastructures), Work Package 4 (Resilience of Networks, Structural Health Monitoring and Asset Management).
@@ Acknowledgements
The improved Loewner Framework implementation (Data file 1) used in this work is openly available from the Zenodo repository at [https://zenodo.org/records/13863292] and the identification results and numerical dataset (Data file 2) supporting this study are openly available from the Zenodo repository at [https://zenodo.org/records/13254981]. This project contains the following underlying data:
@@
Data file 2. Data supporting: Improved tangential interpolation-based multi-input multi-output modal analysis of a full aircraft